# Elastic Constants and Bending Rigidities from Long-Wavelength Perturbation Expansions

### v1

@@ II.2 Long-wavelength acoustic vibrations
where $M_{\kappa}$ is the atomic mass of the $\kappa$-th atom within the unit cell and $\Phi_{\kappa\alpha,\kappa^{\prime}\beta}(\mathbf{q})$ is the so-called dynamical matrix. Such a lattice-dynamical problem is solved by diagonalizing the mass-scaled dynamical matrix $\Phi_{\kappa\alpha,\kappa^{\prime}\beta}(\mathbf{q})/\sqrt{M_{\kappa}M_{\kappa^{\prime}}}$. If there is $n_{\mathrm{a}}$ atoms in the unit cell, one can obtain $3n_{\mathrm{a}}$ phonon modes labeled by the momentum $\mathbf{q}$ and mode index $\nu$, with the corresponding phonon frequencies $\omega_{\nu}(\mathbf{q})$ and mass-weighted eigenvectors $\sqrt{M_{\kappa}}U_{\nu,\kappa\alpha}(\mathbf{q})$. Hereafter, the band index $\nu$ is dropped for clarity. With the knowledge of the second-order interatomic force constants (IFCs) in real space, the dynamical matrix at an arbitrary $\mathbf{q}$ in reciprocal space can be obtained via a Fourier transform:
@@
where $\bm{\tau}^{l}_{\kappa\kappa^{\prime}}\equiv\bm{\tau}_{l\kappa^{\prime}}-\bm{\tau}_{0\kappa}$ is introduced with $\bm{\tau}_{l\kappa}$ denoting the equilibrium position of the $\kappa$-th atom within the unit cell $\mathbf{R}_{l}$. Thanks to the translational invariance, the second-order IFCs only depend on the relative position between the two atoms. As a result, it is sufficient to consider only the second-order IFCs $\Phi^{l}_{\kappa\alpha,\kappa^{\prime}\beta}$ that involve the atom $\kappa^{\prime}$ in the unit cell $l$ and the other atom $\kappa$ in the reference unit cell 00. We note that the real-space IFCs discussed here include macroscopic long-range effects that can be removed to obtain the short-circuit elastic constants, see Sec. II.4. In addition, these IFCs need to satisfy the translational and rotational invariances and the equilibrium conditions for vanishing external stress, as summarized in App. A.
@@
which represent the zeroth- to second-order long-wavelength equations, respectively. These are a set of coupled equations with the perturbation series of phonon frequencies $\omega(\mathbf{q})$ and eigenvectors $U_{\kappa\alpha}(\mathbf{q})$ as the unknowns, which must be solved iteratively [65]. The derivation of their solutions is presented in App. B.
@@
where we have used the definitions given by Eq. (94), (96) and (97) for $\bm{\Gamma}$, $\bm{\Lambda}$ and $\bm{\Upsilon}$ which act as the inverse of the zone-center IFC matrix, internal-strain force-response tensor and internal-strain displacement-response tensor, respectively. The antisymmetry relation $\Phi^{(1),\gamma}_{\kappa\alpha,\kappa^{\prime}\beta}=-\Phi^{(1),\gamma}_{\kappa^{\prime}\beta,\kappa\alpha}$ is also used to obtain Eq. (31). In addition, $T^{\mathrm{CI}}_{\alpha\beta,\gamma\delta}$ and $T^{\rm LM}_{\alpha\gamma,\beta\delta}$ follow the symmetry relations:
@@
$[\omega^{(1)}(\mathbf{q})]^{2}=q_{\gamma}q_{\delta}\frac{T_{\alpha\beta,\gamma\delta}}{M}\overline{U}_{\alpha}\overline{U}_{\beta},$ (34)
@@
with the phase velocity $v=\omega^{(1)}(\mathbf{q})/|\mathbf{q}|$. Such an eigenvalue equation is central to the theory of elastic waves in crystals and can be solved by diagonalizing the positive-definite tensor $M^{-1}\hat{q}_{\gamma}\hat{q}_{\delta}T_{\alpha\beta,\gamma\delta}$, which gives rise to the sound propagation velocity in crystals and the corresponding displacement vector along the $\hat{\mathbf{q}}$ direction. Finally, we show in App. C that the coupled long-wavelength Eqs. (25)–(27) are directly linked to the homogeneous elastic deformation.
@@ II.3 Huang formula for elastic stiffness tensor
with $\rho=M/\Omega$, which implies that the following identity holds
@@
This degree of freedom in defining the elastic tensor stems from using a reciprocal-space formulation as in Eq. (10) or (28) from the lattice-dynamical theory, which is often referred as dynamic elastic tensor in contrast to the static one defined in real space [94, 95]. When considering the symmetry relations of $C_{\alpha\gamma,\beta\delta}$, $T^{\mathrm{CI}}_{\alpha\beta,\gamma\delta}$ and $T^{\rm LM}_{\alpha\gamma,\beta\delta}$ given by Eqs. (3), (32) and (33), respectively, the symmetric elastic tensor is defined as [65, 86]
@@
which is the Huang formula for the elastic stiffness tensor and is the only solution to Eq. (38). It is important to emphasize that the elastic tensor just defined is under short-circuit boundary conditions, where the macroscopic electrostatics must be correctly handled (see Sec. II.4 for details). Thus, the elastic constants of crystals can be fully determined from the two tensors, $T^{\mathrm{CI}}_{\alpha\beta,\gamma\delta}$ and $T^{\rm LM}_{\alpha\gamma,\beta\delta}$, and Eq. (40) is one of the main results of this work.
Due to mechanical stability, the results of $T^{\mathrm{CI}}_{\alpha\beta,\gamma\delta}+T^{\mathrm{CI}}_{\beta\gamma,\alpha\delta}-T^{\mathrm{CI}}_{\beta\delta,\alpha\gamma}$ must be a positive-definite matrix which always increases the total energy of crystals under elastic deformations. By contrast, $T^{\rm LM}_{\alpha\gamma,\beta\delta}$ is in general negative as the relaxation of internal ionic degrees of freedom lowers the total energy, if there is more than one atom in the unit cell without enough site symmetries (e.g. inversion). We note that the long-wavelength equation for acoustic modes and the macroscopic equation of motion for elastic waves are equivalent only if the system is initially free of stress such that the equilibrium conditions $T^{\mathrm{CI}}_{\alpha\beta,\gamma\delta}=T^{\mathrm{CI}}_{\gamma\delta,\alpha\beta}$ hold [see Eq. (85)]. Nevertheless, the equilibrium conditions for vanishing external stress can still hold under isotropic and hydrostatic pressures, $\sigma_{\alpha\beta}=-p\delta_{\alpha\beta}$, where $p$ is a constant. This means that the Huang formula for elastic constants can be still applied to crystals under any symmetry-preserving hydrostatic pressure.
@@ II.4 Separation of long-range electrostatic interactions
$\Phi^{\mathcal{L}}(\mathbf{q})=\lim_{\mathbf{q}\to\mathbf{0}}4\pi\Omega\frac{\ket{\overline{\rho}(\mathbf{q})}\bra{\overline{\rho}(\mathbf{q})}}{\xi(\mathbf{q})},$ (42)
where |$\ket{\overline{\rho}(\mathbf{q})}$ is the unscreened charge density response to a phonon perturbation $\mathbf{q}$ after removing the macroscopic potential, and $\xi(\mathbf{q})$ is the dielectric screening function. They are both analytic in the long-wavelength limit ($\mathbf{q}\to\mathbf{0}$), which can be expanded as a Taylor series:
@@
In practice, to numerically compute the multipole and dielectric tensors, it is advantageous to rewrite Eqs. (43) and (II.4) into a cluster expansion form with a least-square fit to obtain the irreducible tensor components. This rewriting using a cluster expansion model is detailed in App. D and is the method we used in this work. Then, by injecting Eqs. (43) and (II.4) into Eq. (42), we have to $\mathcal{O}(q^{2})$ [93, 88]:
@@
$\Phi^{\rm{DD}}(\mathbf{q})$ $=\frac{4\pi e^{2}}{\Omega}\frac{\ket{Z(\mathbf{q})}\bra{Z(\mathbf{q})}}{\mathbf{q}\cdot\bm{\epsilon}^{\infty}\cdot\mathbf{q}},$ (46)
$\Phi^{\rm{DQ}}(\mathbf{q})$ $=-i\frac{4\pi e^{2}}{2\Omega}\frac{\ket{Z(\mathbf{q})}\bra{Q(\mathbf{q})}-\ket{Q(\mathbf{q})}\bra{Z(\mathbf{q})}}{\mathbf{q}\cdot\bm{\epsilon}^{\infty}\cdot\mathbf{q}},$ (47)
$\Phi^{\rm{QQ}}(\mathbf{q})$ $=\frac{4\pi e^{2}}{4\Omega}\frac{\ket{Q(\mathbf{q})}\bra{Q(\mathbf{q})}}{\mathbf{q}\cdot\bm{\epsilon}^{\infty}\cdot\mathbf{q}},$ (48)
$\Phi^{\rm{DO}}(\mathbf{q})$ $=-\frac{4\pi e^{2}}{6\Omega}\frac{\ket{Z(\mathbf{q})}\bra{O(\mathbf{q})}+\ket{O(\mathbf{q})}\bra{Z(\mathbf{q})}}{\mathbf{q}\cdot\bm{\epsilon}^{\infty}\cdot\mathbf{q}},$ (49)
$\Phi^{\rm{D}\epsilon\rm{D}}(\mathbf{q})$ $=\frac{-4\pi e^{2}}{\Omega}\frac{\ket{Z(\mathbf{q})}\bm{\epsilon}^{(4)}\cdot(\mathbf{q}\!\otimes\!\mathbf{q}\otimes\mathbf{q}\otimes\mathbf{q})\bra{Z(\mathbf{q})}}{(\mathbf{q}\cdot\bm{\epsilon}^{\infty}\cdot\mathbf{q})^{2}},$ (50)
where the superscripts D, Q and O denote the dynamical dipoles, quadrupoles and octupoles, respectively and we have introduced the notations:
$\braket{\kappa\alpha}{Z(\mathbf{q})}$ $=q_{\beta}Z^{\beta}_{\kappa\alpha},$ (51)
$\braket{\kappa\alpha}{Q(\mathbf{q})}$ $=q_{\beta}q_{\gamma}Q^{\beta\gamma}_{\kappa\alpha},$ (52)
$\braket{\kappa\alpha}{O(\mathbf{q})}$ $=q_{\beta}q_{\gamma}q_{\delta}O^{\beta\gamma\delta}_{\kappa\alpha}.$ (53)
Specifically, $\Phi^{\rm{D}\epsilon\rm{D}}(\mathbf{q})$ in Eq. (50) represents the dipole-dipole interaction mediated by dielectric spacial dispersion [93, 88]. For practical calculations of long-range dynamical matrix, the Ewald summation technique as described in Ref. [60] is adopted, and the multipolar electrostatic contributions to dynamical matrix can be rewritten into a summation over reciprocal lattice $\mathbf{G}$ vectors as
$C^{\mathcal{L}}_{\kappa\alpha,\kappa^{\prime}\beta}(\mathbf{q})=$ $\sum_{\mathbf{G}\neq\mathbf{-q}}\braket{\kappa\alpha}{\Phi^{\mathcal{L}}(\mathbf{q+G})}{\kappa^{\prime}\beta}\times$
@@
[…]
where $\|\cdots\|_{1}$ represents the $\ell_{1}$-norm (or Manhanttan distance) and the asterisk means the self-interaction terms are excluded. We use the optimal $\Lambda$ that minimizes Eq. (56) to calculate the long-range part of dynamical matrix and interpolate phonon dispersions shown in Sec. IV, whose value depends on the material of interest and the included multipolar terms. However, in the case of elastic tensor calculations using short-circuit boundary conditions, $\Lambda$ needs to be small such that it only removes the macroscopic $\mathbf{G=0}$ component of the long-range Coulomb interaction. The optimal $\Lambda$ obtained from Eq. (56) generally will be too large, which removes a certain portion of short-range components of Coulomb interactions as well. Therefore, to obtain the short-circuit elastic tensor we minimize the spread of the IFCs in Eq. (56) with the only $\mathbf{G=0}$ taken into account to find the optimal value of $\Lambda$. We find and show in Sec. IV that the elastic tensor is not very sensitive to the choice of the range separation parameters.
@@
Before going further, we would like to stress the significance of the developed long-wavelength perturbation framework. What we have achieved is an expression for the elastic constants which only requires the knowledge of the IFCs in real space without relying on finite differences. This, by itself, has many advantages even if elastic constants are routinely computed by many ab initio codes. However, the benefit of our formulation is that it also gives access to other quantities that are otherwise almost impossible to compute, one such quantity being the bending rigidity tensor.
@@ III Extension to bending rigidities
We begin the study of flexural properties of 2D materials by briefly reviewing the classical Kirchhoff plate theory [106, 107, 108]. For a plate of thickness $h$, it is common to define an auxiliary mid-plane that bisects the plate into two equal halves (i.e. $-h/2\leq z\leq h/2$), and we shall only study the case of such thin plate under small deflections. The Kirchhoff plate theory relies on three assumptions known as Kirchhoff hypotheses [106, 107, 108]: (i) the deflection is small compared to the thickness of plate and the mid-plane remains unstrained; (ii) the plane cross-section remains normal to the mid-plane and to the surfaces before and after bending deformations; and (iii) the plate behaves like linear-elastic materials and the Hooke’s law is valid. Importantly, (ii) implies that the thickness direction is not extensible and the strain components $\varepsilon_{zz}$, $\varepsilon_{xz}$ and $\varepsilon_{yz}$ are vanishing. Consequently, the out-of-plane displacement $w$ becomes independent of the position $z$ as $w(x,y)$, while the in-plane displacement varies linearly in $z$ as $\mathbf{u}^{\parallel}(x,y)=-z\nabla w(x,y)$. For the in-plane stress-strain fields, the constitutive relation described by the generalized Hooke’s law and the corresponding equations of motion for longitudinal and transverse elastic waves in Sec. II.1 still hold for 2D materials, but with the additional flexural contributions to strain tensor. According to the Kirchhoff plate theory, the flexural in-plane strain tensor is given by $\varepsilon_{\alpha\beta}=-z\kappa_{\alpha\beta}$, where $\kappa_{\alpha\beta}=\partial^{2}w/(\partial\tau_{\alpha}\partial\tau_{\beta})$ is the curvature tensor (which should not be confused with the one for atom sites in the unit cell). We therefore focus here on the out-of-plane displacements $w$, bending waves $\omega_{\mathrm{ZA}}(\mathbf{q})$ and their equations of motion.

### v2

@@ II.2 Long-wavelength acoustic vibrations
[…]
@@
[…]
@@
which represent the zeroth- to second-order long-wavelength equations, respectively. These are a set of coupled equations with the perturbation series of phonon frequencies $\omega(\mathbf{q})$ and eigenvectors $U_{\kappa\alpha}(\mathbf{q})$ as the unknowns, which must be solved iteratively [70]. The derivation of their solutions is presented in Appendix B.
@@
where we have used the definitions given by Eq. (94), (96), and (97) for $\bm{\Gamma}$, $\bm{\Lambda}$, and $\bm{\Upsilon}$ which act as the inverse of the zone-center IFC matrix, internal-strain force-response tensor and internal-strain displacement-response tensor, respectively. The antisymmetry relation $\Phi^{(1),\gamma}_{\kappa\alpha,\kappa^{\prime}\beta}=-\Phi^{(1),\gamma}_{\kappa^{\prime}\beta,\kappa\alpha}$ is also used to obtain Eq. (31). In addition, $T^{\mathrm{CI}}_{\alpha\beta,\gamma\delta}$ and $T^{\rm LM}_{\alpha\gamma,\beta\delta}$ follow the symmetry relations:
@@
$\omega^{(1)}(\mathbf{q})]^{2}=q_{\gamma}q_{\delta}\frac{T_{\alpha\beta,\gamma\delta}}{M}\overline{U}_{\alpha}\overline{U}_{\beta},$ (34)
@@
with the phase velocity $v=\omega^{(1)}(\mathbf{q})/|\mathbf{q}|$. Such an eigenvalue equation is central to the theory of elastic waves in crystals and can be solved by diagonalizing the positive-definite tensor $M^{-1}\hat{q}_{\gamma}\hat{q}_{\delta}T_{\alpha\beta,\gamma\delta}$, which gives rise to the sound propagation velocity in crystals and the corresponding displacement vector along the $\hat{\mathbf{q}}$ direction. Finally, we show in Appendix C that the coupled long-wavelength Eqs. (25)–(27) are directly linked to the homogeneous elastic deformation.
@@ II.3 Huang formula for elastic stiffness tensor
with $\rho=M/\Omega$, which implies that the following identity holds:
@@
This degree of freedom in defining the elastic tensor stems from using a reciprocal-space formulation as in Eq. (10) or (28) from the lattice-dynamical theory, which is often referred to as dynamic elastic tensor in contrast to the static one defined in real space [101, 102]. When considering the symmetry relations of $C_{\alpha\gamma,\beta\delta}$, $T^{\mathrm{CI}}_{\alpha\beta,\gamma\delta}$ and $T^{\rm LM}_{\alpha\gamma,\beta\delta}$ given by Eqs. (3), (32), and (33), respectively, the symmetric elastic tensor is defined as [70, 92]
@@
which is the Huang formula for the elastic stiffness tensor and is the only solution to Eq. (38). It is important to emphasize that the elastic tensor just defined is under short-circuit boundary conditions, where the macroscopic electrostatics must be correctly handled (see Sec. II.4 for details). Thus, the elastic constants of crystals can be fully determined from the two tensors, $T^{\mathrm{CI}}_{\alpha\beta,\gamma\delta}$ and $T^{\rm LM}_{\alpha\gamma,\beta\delta}$.
Due to mechanical stability, the results of $T^{\mathrm{CI}}_{\alpha\beta,\gamma\delta}+T^{\mathrm{CI}}_{\beta\gamma,\alpha\delta}-T^{\mathrm{CI}}_{\beta\delta,\alpha\gamma}$ must be a positive-definite matrix, which always increases the total energy of crystals under elastic deformations. By contrast, $T^{\rm LM}_{\alpha\gamma,\beta\delta}$ is in general negative as the relaxation of internal ionic degrees of freedom lowers the total energy, if there is more than one atom in the unit cell without enough site symmetries (e.g. inversion). We note that the long-wavelength equation for acoustic modes and the macroscopic equation of motion for elastic waves are equivalent only if the system is initially free of stress such that the equilibrium conditions $T^{\mathrm{CI}}_{\alpha\beta,\gamma\delta}=T^{\mathrm{CI}}_{\gamma\delta,\alpha\beta}$ hold [see Eq. (85)]. Nevertheless, the equilibrium conditions for vanishing external stress can still hold under isotropic and hydrostatic pressures, $\sigma_{\alpha\beta}=-p\delta_{\alpha\beta}$, where $p$ is a constant. This means that the Huang formula for elastic constants can be still applied to crystals under any symmetry-preserving hydrostatic pressure.
@@ II.4 Separation of long-range electrostatic interactions
[…]
@@
In practice, to numerically compute the multipole and dielectric tensors, it is advantageous to rewrite Eqs. (43) and (II.4) into a cluster expansion form with a least-squares fit to obtain the irreducible tensor components. This rewriting using a cluster expansion model is detailed in Appendix D and is the method we used in this work. Then, by injecting Eqs. (43) and (II.4) into Eq. (42), we have to $\mathcal{O}(q^{2})$ [100, 95]:
@@
[…]
where the superscripts D, Q, and O denote the dynamical dipoles, quadrupoles and octupoles, respectively, and we have introduced the notations:
$\braket{\kappa\alpha|Z(\mathbf{q})}$ $=q_{\beta}Z^{\beta}_{\kappa\alpha},$ (51)
$\braket{\kappa\alpha|Q(\mathbf{q})}$ $=q_{\beta}q_{\gamma}Q^{\beta\gamma}_{\kappa\alpha},$ (52)
$\braket{\kappa\alpha|O(\mathbf{q})}$ $=q_{\beta}q_{\gamma}q_{\delta}O^{\beta\gamma\delta}_{\kappa\alpha}.$ (53)
Specifically, $\Phi^{\rm{D}\epsilon\rm{D}}(\mathbf{q})$ in Eq. (50) represents the dipole-dipole interaction mediated by dielectric spacial dispersion [100, 95]. For practical calculations of the long-range dynamical matrix, the Ewald summation technique as described in Ref. [64] is adopted, and the multipolar electrostatic contributions to dynamical matrix can be rewritten into a summation over reciprocal lattice $\mathbf{G}$ vectors as
$C^{\mathcal{L}}_{\kappa\alpha,\kappa^{\prime}\beta}(\mathbf{q})=$ $\sum_{\mathbf{G}\neq\mathbf{-q}}\braket{\kappa\alpha|\Phi^{\mathcal{L}}(\mathbf{q+G})|\kappa^{\prime}\beta}\times$
@@
$\Phi^{\mathcal{S}}(\mathbf{R}_{l})=\overunderset{*}{\Lambda}{\arg\min}\big\|\Phi^{\mathcal{S}}(\mathbf{R}_{l})\big\|_{1},$ (56)
where $\|\cdots\|_{1}$ represents the $\ell_{1}$-norm (or Manhanttan distance) and the asterisk means the self-interaction terms are excluded. We use the optimal $\Lambda$ that minimizes Eq. (56) to calculate the long-range part of dynamical matrix and interpolate phonon dispersions shown in Sec. IV, whose value depends on the material of interest and the included multipolar terms. However, in the case of elastic tensor calculations using short-circuit boundary conditions, $\Lambda$ needs to be small such that it only removes the macroscopic $\mathbf{G=0}$ component of the long-range Coulomb interaction. The optimal $\Lambda$ obtained from Eq. (56) generally will be too large, which removes a certain portion of short-range components of Coulomb interactions as well. Therefore, to obtain the short-circuit elastic tensor we minimize the sum of short-range IFCs in Eq. (56) with the only $\mathbf{G=0}$ taken into account to find the optimal value of $\Lambda$. We find and show in Sec. IV that the elastic tensor is not very sensitive to the choice of the range separation parameters. The removal of such $\mathbf{G=0}$ component is also consistent with the DFPT formulation of elastic tensor, which excludes the macroscopic Hartree, local pseudopotential, and Ewald terms [100].
@@
Before going further, we would like to stress the significance of the developed long-wavelength perturbation framework. What we have achieved is an expression for the elastic constants, which only requires the knowledge of IFCs in real space, avoiding the finite difference of stress tensor, and IFCs can be obtained via DFPT [64, 66] or small displacement method [113], making it amenable to any atomistic code that provides forces. This, by itself, has many advantages even if elastic constants are routinely computed by many ab initio codes. In particular, the Huang formula is independent on the PBCs, which are implicitly treated by constructing the optimal Wigner-Seitz supercell in lattice dynamical calculations [97], and this means Eq. (40) can be applied equivalently to any dimension of systems. However, the benefit of our formulation is that it also gives access to other quantities that are otherwise almost impossible to compute, one such quantity being the bending rigidity tensor. Furthermore, the extracted high-order multipole tensors from our generalized cluster expansion approach will enable the study of piezoelectricity [114] and flexoelectricity [100], and provide an accurate description of long-range electron-phonon coupling beyond Fröhlich interactions.
@@ III Extension to bending rigidities
We begin the study of flexural properties of 2D materials by briefly reviewing the classical Kirchhoff plate theory [115, 116, 117]. For a plate of thickness $h$, it is common to define an auxiliary midplane that bisects the plate into two equal halves (i.e. $h/2\leq z\leq h/2$), and we shall only study the case of such thin plate under small deflections. The Kirchhoff plate theory relies on three assumptions known as Kirchhoff hypotheses [115, 116, 117]: (i) the deflection is small compared to the thickness of plate and the midplane remains unstrained; (ii) the plane cross section remains normal to the midplane and to the surfaces before and after bending deformations; and (iii) the plate behaves like linear-elastic materials and the Hooke’s law is valid. Importantly, (ii) implies that the thickness direction is not extensible and the strain components $\varepsilon_{zz}$, $\varepsilon_{xz}$ and $\varepsilon_{yz}$ are vanishing. Consequently, the out-of-plane displacement $w$ becomes independent of the position $z$ as $w(x,y)$, while the in-plane displacement varies linearly in $z$ as $\mathbf{u}^{\parallel}(x,y)=-z\nabla w(x,y)$. For the in-plane stress-strain fields, the constitutive relation described by the generalized Hooke’s law and the corresponding equations of motion for longitudinal and transverse elastic waves in Sec. II.1 still hold for 2D materials, but with the additional flexural contributions to strain tensor. According to the Kirchhoff plate theory, the flexural in-plane strain tensor is given by $\varepsilon_{\alpha\beta}=-z\kappa_{\alpha\beta}$, where $\kappa_{\alpha\beta}=\partial^{2}w/(\partial\tau_{\alpha}\partial\tau_{\beta})$ is the curvature tensor (which should not be confused with the one for atom sites in the unit cell). We therefore focus here on the out-of-plane displacements $w$, bending waves $\omega_{\mathrm{ZA}}(\mathbf{q})$, and their equations of motion.